OpenStudy (anonymous):

if a plane can travel 490 miles per hour with the wind and 410 miles per hour against the wind find the speed of the plane with out wind and with the speed of the wind?

OpenStudy (inkyvoyd):

okay so here you have 2 speeds you do not know: the speed of a plane and the speed of the wind. Defining the speed of the plane to be positive, if the plane travels with the wind then x+y=490 If the plane travels against the wind then x-y=410 Do you know how to solve this system?

OpenStudy (anonymous):

yes

OpenStudy (inkyvoyd):

okay, let me know what you get :)

OpenStudy (anonymous):

wait im a little confused to do i add these equations to get my answer?

OpenStudy (inkyvoyd):

that' one way to do it x+y=490 x-y=410 +------- 2x+0y=0-- so you get 2x+0y=900

OpenStudy (anonymous):

so x=450

OpenStudy (inkyvoyd):

yes - now do you know how to find the speed of the wind?

OpenStudy (anonymous):

negative

OpenStudy (inkyvoyd):

just put it in the original equation. x+y=490 we know that x=450 so now what?

OpenStudy (anonymous):

k. got it y=40.

OpenStudy (inkyvoyd):

yup

OpenStudy (anonymous):

thanksfor your help

OpenStudy (inkyvoyd):

np :)